If f(x)=x^3 & g(x)=x+5, which is equivalent form of f(g(x))?
Answer by jim_thompson5910(35256)

f(x)=x^3

f(g(x))=(g(x))^3

f(g(x))=(x+5)^3


So f(g(x))=(x+5)^3